# Theodore S. Rappaport, Mingjun Ying, Nicola Piovesan, Antonio De Domenico, and Dipankar Shakya, “Waste Factor and Waste Figure: A Unified

### v1

@@ IV-G Analyzing Quiescent Power in Amplifiers
Incorporating quiescent power into $P_{\text{non-signal}}$ would be misleading, as it would suggest null total power usage in the absence of an output signal, which contradicts the actual power dynamics. Consequently, it is imperative to treat the quiescent or static/stand-by power of a PA as non-path power to ensure an accurate assessment of a system’s overall power consumption along an active signal path.
IV-H Waste Factor of the Wireless Channel
@@
$=\frac{1}{G_{\text{C}}}=L_{\text{C}},$ (70)
@@
$L_{\text{C}}^{\text{eff}}$ $=\frac{L_{\text{C}}}{G_{\text{TX}}^{\text{ant}}G_{\text{RX}}^{\text{ant}}}$
@@ IV-I Waste Factor of RAN
$=W_{\text{Ant}}+\frac{W_{\text{PA}}-1}{G_{\text{Ant}}}+\frac{W_{\text{PS}}-1}{G_{\text{PA}}G_{\text{Ant}}}$ (73)
@@
$W_{0}^{\textrm{RU}}=W_{\textrm{Mix}}+\frac{W_{\textrm{DAC}}-1}{G_{\textrm{Mix}}}$ (74)
@@
$+\frac{W_{\text{Mix}}-1}{G_{\text{PS}}G_{\text{PA}}G_{\text{Ant}}}+\frac{W_{\text{DAC}}-1}{G_{\text{PS}}G_{\text{PA}}G_{\text{Ant}}G_{\text{Mix}}}$ (75)
@@
$=W_{\text{PS}}+\frac{W_{\text{LNA}}-1}{G_{\text{PS}}}+\frac{W_{\text{Ant}}-1}{G_{\text{PS}}G_{\text{LNA}}}$ (76)
@@
$=W_{\text{Mix}}+\frac{W_{\text{PS}}-1}{G_{\text{Mix}}}+\frac{W_{\text{LNA}}-1}{G_{\text{Mix}}G_{\text{PS}}}+\frac{W_{\text{Ant}}-1}{G_{\text{Mix}}G_{\text{PS}}G_{\text{LNA}}}$ (78)
@@
IV-J Computing W through measurements
@@
$P_{\text{consumed,path}}=P_{\text{consumed,total}}-P_{\text{non-path}}$ (80)
@@
$W_{\text{RU}}=\frac{P_{\text{consumed,path}}}{P_{\text{signal}}}$ (81)
@@ V-A Impact of channel and components characteristics on RAN power waste
of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.60606060of the Communications Society (OJ-COMS), May 2024, pp. 1–26.70707070of the Communications Society (OJ-COMS), May 2024, pp. 1–26.80808080of the Communications Society (OJ-COMS), May 2024, pp. 1–26.90909090of the Communications Society (OJ-COMS), May 2024, pp. 1–26.100100100100of the Communications Society (OJ-COMS), May 2024, pp. 1–26.60606060of the Communications Society (OJ-COMS), May 2024, pp. 1–26.70707070of the Communications Society (OJ-COMS), May 2024, pp. 1–26.80808080of the Communications Society (OJ-COMS), May 2024, pp. 1–26.90909090of the Communications Society (OJ-COMS), May 2024, pp. 1–26.100100100100of the Communications Society (OJ-COMS), May 2024, pp. 1–26.110110110110of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.3dB gainof the Communications Society (OJ-COMS), May 2024, pp. 1–26.$WF_{\textrm{C}}$ [dB]of the Communications Society (OJ-COMS), May 2024, pp. 1–26.$WF_{\textrm{System}}$ [dB]of the Communications Society (OJ-COMS), May 2024, pp. 1–26.baselinereduced $W_{\textrm{RU}}$reduced $W_{\textrm{ UE}}$improved $G_{\textrm{UE}}$
@@
This section presents a comprehensive quantitative evaluation of W of the system that encompasses the RU, the wireless communication channel, and the UE. Our methodology for evaluating the waste factor integrates a detailed analysis of the constituent components within both the RU and UE, building upon the foundational frameworks outlined in Section IV-IV-E and Section IV-IV-F (e.g. the mathematical formulations presented in eq. (75) and eq. (78) to compute W for the RU and UE, respectively).
@@ V-B Comparison of RU power waste using W
In this section, we assess W of two distinct RUs through measurements, as shown in Section IV-IV-J.
@@
When adopting the energy efficiency metric for the RU defined by ITU in [22] and presented in Section IV-IV-C (see eq. 41) we obtain the same energy efficiency values for both RUs
@@
V-C W to drive decision making
W, when coupled with other relevant network key performance indicators, can serve as a powerful tool to drive decision-making in the design of a strategy to improve the energy efficiency of any communication network.
of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26. Low of the Communications Society (OJ-COMS), May 2024, pp. 1–26.Lowof the Communications Society (OJ-COMS), May 2024, pp. 1–26. Data rate of the Communications Society (OJ-COMS), May 2024, pp. 1–26.Waste Factorof the Communications Society (OJ-COMS), May 2024, pp. 1–26. High of the Communications Society (OJ-COMS), May 2024, pp. 1–26.High of the Communications Society (OJ-COMS), May 2024, pp. 1–26. Optimal of the Communications Society (OJ-COMS), May 2024, pp. 1–26. Optimize scheduled transmit power of the Communications Society (OJ-COMS), May 2024, pp. 1–26. Deploy efficient hardware, smaller cells of the Communications Society (OJ-COMS), May 2024, pp. 1–26. Increase transmit power, bandwidth, carrier aggregation
@@ V-C W to drive decision making
of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26.of the Communications Society (OJ-COMS), May 2024, pp. 1–26. Low of the Communications Society (OJ-COMS), May 2024, pp. 1–26.Lowof the Communications Society (OJ-COMS), May 2024, pp. 1–26. Consumed
@@
of the Communications Society (OJ-COMS), May 2024, pp. 1–26.Waste Factorof the Communications Society (OJ-COMS), May 2024, pp. 1–26. High of the Communications Society (OJ-COMS), May 2024, pp. 1–26.High of the Communications Society (OJ-COMS), May 2024, pp. 1–26. Shutdown methods, efficient cooling of the Communications Society (OJ-COMS), May 2024, pp. 1–26. Optimize scheduled transmit power of the Communications Society (OJ-COMS), May 2024, pp. 1–26. Deploy efficient hardware of the Communications Society (OJ-COMS), May 2024, pp. 1–26. Optimal
@@ V-D Assessing the energy efficiency of distributed MU-MIMO systems using W
A single Waste Factor value for the entire distributed MU-MIMO system is calculated following the methodology outlined in Section V-III-C. In Fig. 16, the BSs shown within the leftmost dashed box represent the M (range from 1 to 20) BSs used in the simulation, while the UEs shown within the middle dotted box represent the N UEs (N = 1024), and each UE is equipped with a single antenna (note: using eqs. (30) and (34) in Section III, we could easily consider each of the UE together with the BSs in Fig. 16 being a MIMO system).
@@
To determine the W and total power consumption of the entire distributed MU-MIMO network composed of M BSs and 1024 UEs, we first determine the W for each BS and UE separately using eqs. (75) and (78), respectively. Based on the detailed gain and efficiency and DC power consumption from data sheets of real-world components, we apply equations (IV-E) and (56) for passive and active components in the RAN system given in Section IV, and (II-C) to compute the Waste Factor for a UE and BS, and we assume all BSs have identical $W_{\text{BS}}=15$ and $G_{\text{BS}}=30\,\text{dB}$, and that all 1024 UEs have the same $W_{\text{UE}}=33$ and $G_{\text{UE}}=11\,\text{dB}$. Also, we assume that the non-path power for each BS is $P_{\text{non-path}}^{\text{BS}}=140\,\text{W}$ (consistent with the values used in Section V), and the non-path power for each UE is $P_{\text{non-path}}^{\text{UE}}=1\,\text{W}$, which is a typical value since the display and the processor of the UE will averagely consume around 1 W [75]. All channels shown in Fig. 16 are different and yield different values of $W_{\text{C}}$ based on the simulation of the propagation, but we dynamically adjust the transmit power in the simulation (to model distributed MU-MIMO) for each of the parallel BSs to achieve a 10 dB SNR at the input of all UEs, however, in some cases, unrealistic propagation conditions for omnidirectional antennas do not allow all UE to achieve a 10 dB SNR as the BSs transmit power is limited to 10 dBm for each BS-UE link.
@@
$W_{\text{N}||}=\frac{\sum_{n=1}^{1024}(P_{2,n}W_{\text{UE}})}{\sum_{n=1}^{1024}P_{2,n}}=W_{\text{UE}}.$ (87)
@@
Fig. 19 illustrates the total power consumption of a distributed MU-MIMO system utilizing omnidirectional antennas across three carrier frequencies. By aggregating signal power from 1024 UEs at an imagined sink, and employing eq. (90), the simulation calculates the total power consumption of the simulated system. The total power consumption at 3.5 GHz is considerably lower compared to higher frequencies due to the nonsensical assumption that omnidirectional antennas are used in all three bands such that channel loss increases by the square of frequency without any benefit of channel gain that comes from directional antennas. Here, we see clearly how the spherical radiation spreading for omnidirectional antennas [33, 76, 77, 74] include more propagation loss at 17 GHz, where the total power consumption is slightly lower than at 28 GHz, but higher than 3.5 GHz. At 28 GHz, the total power consumption is the highest among all three frequencies and increases with the number of BSs and is 30% higher than at 3.5 GHz with 20 BSs.
Fig. 20 illustrates a more realistic mobile scenario where $\textit{WF}_{\text{System}}$ is computed for a distributed MU-MIMO system using directional antennas. Here, the $\textit{WF}_{\text{System}}$ is remarkably reduced for all three bands as compared to Fig. 18, demonstrating how antenna gains yield less lossy channels that offer dramatic power savings and improved energy efficiency as compared to the omnidirectional systems. $\textit{WF}_{\text{System}}$ at 3.5 GHz, though reduced compared to the omnidirectional scenario of Fig. 18, is larger than the other higher frequencies due to the limited antenna gains at this band (see Table V). At 17 GHz, the $\textit{WF}_{\text{System}}$ decreases rapidly with up to 5 BSs from 81 dB with 1 BS to 79 dB with 5 BSs, but the reduction rate stabilizes as the BS count rises further, highlighting the efficiency gains from directional antennas initially. For the 28 GHz band, the $\textit{WF}_{\text{System}}$ drops significantly with increasing BS density, dropping from 78 dB for 1 BS to 76 dB for 5 BSs and 75 dB for 15 BSs, showing a clear advantage due to directional antennas which counter the propagation and shadowing loss effects. The $\textit{WF}_{\text{System}}$ with 20 BSs is 12 dB lower than at 3.5 GHz with 20 BSs, highlighting the substantial power efficiency gains (more than an order of magnitude of power savings) at higher frequencies. Even greater efficiencies occur due to the fact that 5G mmWave technology is inherently more power efficient (by 2$\times$) than legacy 4G technology (see Section IV), thereby showing the move to mmWave small cells offers much greater power efficiency over today’s cellular networks.
@@
VI Conclusion

### v2

@@ IV-F Analyzing Quiescent Power in Amplifiers
Incorporating quiescent power into $P_{\text{non-signal}}$ would be misleading, as it would suggest null total power usage in the absence of an output signal, which contradicts the actual power dynamics. Consequently, it is imperative to treat the quiescent or static/stand-by power of a PA as non-path power to ensure an accurate assessment of a system overall power consumption along an active signal path. Our analysis on LNA and PA further illustrates how W is applicable to a wide range of transmission schemes, including single carrier, OFDM, and carrier aggregation.
IV-G Waste Factor of the Wireless Channel
@@
$=\frac{1}{G_{\text{C}}}=L_{\text{C}}.$ (70)
@@
$L_{\text{C}}^{\text{eff}}$ $=\frac{L_{\text{C}}}{G_{\text{TX}}^{\text{ant}}G_{\text{RX}}^{\text{ant}}},$
@@ IV-H Waste Factor of RAN
$=W_{\text{Ant}}+\frac{W_{\text{PA}}-1}{G_{\text{Ant}}}+\frac{W_{\text{PS}}-1}{G_{\text{PA}}G_{\text{Ant}}}.$ (73)
@@
$W_{0}^{\textrm{RU}}=W_{\textrm{Mix}}+\frac{W_{\textrm{DAC}}-1}{G_{\textrm{Mix}}}.$ (74)
@@
$+\frac{W_{\text{Mix}}-1}{G_{\text{PS}}G_{\text{PA}}G_{\text{Ant}}}+\frac{W_{\text{DAC}}-1}{G_{\text{PS}}G_{\text{PA}}G_{\text{Ant}}G_{\text{Mix}}}.$ (75)
@@
$=W_{\text{PS}}+\frac{W_{\text{LNA}}-1}{G_{\text{PS}}}+\frac{W_{\text{Ant}}-1}{G_{\text{PS}}G_{\text{LNA}}}.$ (76)
@@
$=W_{\text{Mix}}+\frac{W_{\text{PS}}-1}{G_{\text{Mix}}}+\frac{W_{\text{LNA}}-1}{G_{\text{Mix}}G_{\text{PS}}}+\frac{W_{\text{Ant}}-1}{G_{\text{Mix}}G_{\text{PS}}G_{\text{LNA}}}.$ (78)
@@
IV-I Computing W through measurements
@@
$P_{\text{consumed,path}}=P_{\text{consumed,total}}-P_{\text{non-path}}.$ (80)
@@
$W_{\text{RU}}=\frac{P_{\text{consumed,path}}}{P_{\text{signal}}}.$ (81)
@@ V-A Impact of channel and components characteristics on RAN power waste
[…]
@@
This section presents a comprehensive quantitative evaluation of W of the system that encompasses the RU, the wireless communication channel, and the UE. Our methodology for evaluating the waste factor integrates a detailed analysis of the constituent components within both the RU and UE, building upon the foundational frameworks outlined in Section IV-IV-D and Section IV-IV-E (e.g. the mathematical formulations presented in eq. (75) and eq. (78) to compute W for the RU and UE, respectively).
@@ V-B Comparison of RU power waste using W
In this section, we assess W of two distinct RUs through measurements, as shown in Section IV-IV-I.
@@
When adopting the energy efficiency metric for the RU defined by ITU in [26] and presented in Section IV-IV-B (see eq. (42)) we obtain the same energy efficiency values for both RUs
@@
V-C Waste Factor: A Key Value Indicator for sustainability and optimization
As demonstrated in Section IV-IV-C, W offers a significant advantage over existing standard EE metrics from ITU, ETSI, and 3GPP because W is unaffected by traffic load variations at which it is quantified. This unique characteristic of W through the parsing of signal path and non-signal path energy consumption, and the partitioning of quiescent or static/stand-by power from energy consumed during information flow, enables a more accurate assessment of the efficiency of RUs for data communications. Currently, EE metrics need to be frequently recalculated to evaluate the efficiency of RUs under varying traffic conditions. In contrast, each RU has a unique W determining the intrinsic energy efficiency in data transmission. The nature of W has a substantial impact on network optimization and provides mobile network operators with a means to identify less efficient RUs independent of their operating condition. For instance, these inefficient RUs can be replaced with more energy-efficient hardware or can be shut down during low-traffic periods, with users being transferred to more efficient RUs (e.g., with lower W). Moreover, W, when coupled with other relevant network Key Performance Indicators (KPIs) or Key Value Indicators (KVIs), offers a powerful tool to drive decision-making strategies that improve the energy efficiency of any communication network.
Low Low Data rate Waste Factor High High Optimal Optimize scheduled transmit power Deploy efficient hardware, smaller cells Increase transmit power, bandwidth, carrier aggregation
@@ V-C Waste Factor: A Key Value Indicator for sustainability and optimization
Low Low Consumed
@@
Waste Factor High High Shutdown methods, efficient cooling Optimize scheduled transmit power Deploy efficient hardware Optimal
@@ V-D Assessing the energy efficiency of distributed MU-MIMO systems using W
A single Waste Factor value for the entire distributed MU-MIMO system is calculated following the methodology outlined in Section III-III-C. In Fig. 16, the BSs shown within the leftmost dashed box represent the M (range from 1 to 20) BSs used in the simulation, while the UEs shown within the middle dotted box represent the N UEs (N = 1024), and each UE is equipped with a single antenna (note: using eqs. (31) and (35) in Section III, we could easily consider each of the UE together with the BSs in Fig. 16 being a MIMO system).
@@
To determine the W and total power consumption of the entire distributed MU-MIMO network composed of M BSs and 1024 UEs, we first determine the W for each BS and UE separately using eqs. (75) and (78), respectively. Based on the detailed gain and efficiency and DC power consumption from data sheets of real-world components, we apply equations (IV-D) and (57) for passive and active components in the RAN system given in Section IV, and (II-C) to compute the Waste Factor for a UE and BS, and we assume all BSs have identical $W_{\text{BS}}=15$ and $G_{\text{BS}}=30\,\text{dB}$, and that all 1024 UEs have the same $W_{\text{UE}}=33$ and $G_{\text{UE}}=11\,\text{dB}$. Also, we assume that the non-path power for each BS is $P_{\text{non-path}}^{\text{BS}}=140\,\text{W}$ (consistent with the values used in Section V-B), and the non-path power for each UE is $P_{\text{non-path}}^{\text{UE}}=1\,\text{W}$, which is a typical value since the display and the processor of the UE will averagely consume around 1 W [75]. All channels shown in Fig. 16 are different and yield different values of $W_{\text{C}}$ based on the simulation of the propagation, but we dynamically adjust the transmit power in the simulation (to model distributed MU-MIMO) for each of the parallel BSs to achieve a 10 dB SNR at the input of all UEs, however, in some cases, unrealistic propagation conditions for omnidirectional antennas do not allow all UE to achieve a 10 dB SNR as the BSs transmit power is limited to 10 dBm for each BS-UE link.
@@
$W_{\text{N}||}=\frac{\sum_{n=1}^{1024}(P_{2,n}W_{\text{UE}})}{\sum_{n=1}^{1024}P_{2,n}}=W_{\text{UE}},$ (87)
@@
Fig. 19 illustrates the total power consumption of a distributed MU-MIMO system utilizing omnidirectional antennas across three carrier frequencies. By aggregating signal power from 1024 UEs at an imagined sink, and employing eq. (90), the simulation calculates the total power consumption of the simulated system.333 In the assessment of energy efficiency using W, it is essential to clarify the relationship between power (P) and energy (E). The relationship between these quantities over a unit of time T in seconds is given by E (Joules) = P (Watts) $\times$ T (s). In the context of our study, this implies that power efficiency and energy efficiency can be considered synonymous over a fixed time period. This clarification ensures that our analysis of true power levels over space (W/km2) is well-grounded and accurately interpreted. The total power consumption at 3.5 GHz is considerably lower compared to higher frequencies due to the nonsensical assumption that omnidirectional antennas are used in all three bands such that channel loss increases by the square of frequency without any benefit of channel gain that comes from directional antennas. Here, we see clearly how the spherical radiation spreading for omnidirectional antennas [37, 76, 77, 74] include more propagation loss at 17 GHz, where the total power consumption is slightly lower than at 28 GHz, but higher than 3.5 GHz. At 28 GHz, the total power consumption is the highest among all three frequencies and increases with the number of BSs and is 30% higher than at 3.5 GHz with 20 BSs.
Fig. 20 illustrates a more realistic mobile scenario where $\textit{WF}_{\text{System}}$ is computed for a distributed MU-MIMO system using directional antennas. Here, the $\textit{WF}_{\text{System}}$ is remarkably reduced for all three bands as compared to Fig. 18, demonstrating how antenna gains yield less lossy channels that offer dramatic power savings and improved energy efficiency as compared to the omnidirectional systems. $\textit{WF}_{\text{System}}$ at 3.5 GHz, though reduced compared to the omnidirectional scenario of Fig. 18, is larger than the other higher frequencies due to the limited antenna gains at this band (see Table V). At 17 GHz, the $\textit{WF}_{\text{System}}$ decreases rapidly with up to 5 BSs from 81 dB with 1 BS to 79 dB with 5 BSs, but the reduction rate stabilizes as the BS count rises further, highlighting the efficiency gains from directional antennas initially. For the 28 GHz band, the $\textit{WF}_{\text{System}}$ drops significantly with increasing BS density, dropping from 78 dB for 1 BS to 76 dB for 5 BSs and 75 dB for 15 BSs, showing a clear advantage due to directional antennas which counter the propagation and shadowing loss effects. The $\textit{WF}_{\text{System}}$ with 20 BSs is 12 dB lower than at 3.5 GHz with 20 BSs, highlighting the substantial power efficiency gains (more than an order of magnitude of power savings) at higher frequencies. Even greater efficiencies occur due to the fact that 5G mmWave technology is inherently more power efficient (by 2$\times$) than legacy 4G technology (see Section IV), thereby showing the move to mmWave small cells offers much greater power efficiency over current cellular networks.
@@
  5. 5.
    In analyzing the power consumption across different antenna configurations, our findings in Figure 19 and Figure 21 distinctly highlight the energy efficiency gains of directional antennas. Specifically, for 20 BSs operating at 28 GHz with directional antennas, the power consumption is markedly lower, at only 4.4 kW/km2, compared to 9 kW/km2 with omnidirectional antennas. This trend is consistent across different densities of BSs. Utilizing $W$ simplifies the calculation of total power and energy consumption in real networks, providing a clear and measurable insight into actual power and energy levels.
VI Limitations and Challenges of Waste Factor Adoption
While W presents a novel approach to evaluate energy efficiency in communication systems, some limitations and challenges need to be addressed for its broader adoption and application in the future:
  • Linearity Requirement: One of the primary limitations of W is that it requires the components of the system to exhibit linear behavior. Real-world communication systems include non-linear components, which complicates the accurate application of W. The applicability to linearized models, and the sensitivity of Waste Factor in the face of real-world non-linearities are open areas of work.
  • Novelty and Lack of Comprehensive Application: As W is a relatively new concept, there is limited practical experience in its application. Researchers and engineers are still exploring how to fully utilize W in various contexts. This lack of comprehensive application knowledge means that further studies and practical implementations are necessary to understand its full potential and limitations.
  • Modeling Various Quiescent States: Accurate modeling of different quiescent or “stand-by” states within communication systems may be challenging. The ability to effectively model these states and to assign the power consumption as being on the signal path or off the signal path is crucial for the successful usage of W, and methodologies must be developed and refined through experience, learning, and trial-and-error.
  • Unexplored Network Architectures: Certain network architectures, such as ring networking structures with multiple switches and routers, have not yet been thoroughly analyzed using W. These architectures present unique challenges and opportunities for applying W, and further research is needed to explore these areas.
  • Combining Method Variability: The method of combining signals in the system plays a crucial role in determining the appropriate Waste Factor formula. In some network scenarios, parts of the devices may use non-coherent combining while others use coherent combining. This variability poses an open challenge that needs to be addressed further to ensure accurate and effective application of W.
By acknowledging these limitations and challenges, we aim to provide a balanced perspective on the potential of the Waste Factor and highlight the need for ongoing research and practical experimentation to fully realize its benefits in improving energy efficiency in communication systems.
VII Conclusion